If despositor puts $100 in a bank account that’s earns 4 percent interest compounded annually how much will be in the account after 5 years?
1
Expert's answer
2020-10-16T09:33:42-0400

"A=P(1+\cfrac{r}{100})^{nt}"

P=$100

r=4%

n=1

t=5

"A=100(1+\cfrac{4}{100})^5"


"A=100(1.04)^5"


"A=121.66529024=121.67"

=$121.67